1. thirteen more than four times a number is -91. find the number
2. the total cost of a suit and a coat is $291. the cost of the coat is twice as much as the suit, how much does the coat cost?
3. steave hit 4 more runs than twice the number of runs larry hit. together they hit 10 runs. how many runs did steave hit?

Answer by ankor@dixie-net.com(22740)

Here are the equations for each problem
:
1. thirteen more than four times a number is -91. find the number
4x + 13 = -91
:
2. the total cost of a suit and a coat is $291.
the cost of the coat is twice as much as the suit,
how much does the coat cost?
s + c = 291
c = 2s
Replace c with 2s in the 1st equation, find s, then find c
:
3. steave hit 4 more runs than twice the number of runs larry hit.
together they hit 10 runs.
how many runs did steave hit?
S = 2L + 4
S + L = 10
L = (10-S)
Replace L with (10-S) in the 1st equation, solve for s